OpenStudy (anonymous):

PLEASE HELP Two equations are given below: m + 4n = 8 m = n − 2 What is the solution to the set of equations in the form (m, n)? (4, 6) (2, 4) (0, 2) (6, 8)

Vocaloid (vocaloid):

first, let's leave the second equation alone for now and just look at m+4n = 8 what do you get when you subtract 4n from each side?

OpenStudy (anonymous):

M-N=4

Vocaloid (vocaloid):

hm, not quite, let's just try the left side first what happens when you subtract 4n from m+4n?

OpenStudy (anonymous):

M

Vocaloid (vocaloid):

right! now let's look at the right side what happens when you subtract 4n from 8?

OpenStudy (anonymous):

8N?

Vocaloid (vocaloid):

nope, since 8 has no variable and 4n has a variable, it's just 8 - 4n understand so far?

OpenStudy (anonymous):

Yes

Vocaloid (vocaloid):

ok, so the left side of the equation is m, and the right side of the equation is 8 - 4n, which gives us m = 8 - 4n with me so far?

OpenStudy (anonymous):

yes

Vocaloid (vocaloid):

good, so our first equation is m = 8 - 4n and our second equation is m = n - 2 we have m by itself on the left side in both equations, so we can set the two equations equal to each other 8 - 4n = n - 2 can you try solving for n now?

OpenStudy (anonymous):

I thing so would it be 4,6

Vocaloid (vocaloid):

not quite, let's just take this one step at a time 8 - 4n = n - 2 what happens when you add 2 to each side?

OpenStudy (anonymous):

what do you mean

Vocaloid (vocaloid):

add 2 to the left side, add 2 to the right side, and tell me what is on each side

OpenStudy (anonymous):

10-6n=2n-4

Vocaloid (vocaloid):

not quite, remember we want to leave 4n and n alone because these are variables

OpenStudy (anonymous):

oh

Vocaloid (vocaloid):

left side is 8 - 4n, what do you get when you add 2?

OpenStudy (anonymous):

10-4n

Vocaloid (vocaloid):

great! the right side is n - 2, what do you get when you add 2?

OpenStudy (anonymous):

n-4

Vocaloid (vocaloid):

not quite, what is n-2+2?

OpenStudy (anonymous):

2

Vocaloid (vocaloid):

not quite, think of it this way n-2+2 means start with n, take away 2, then add 2 back what is n-2+2?

OpenStudy (anonymous):

2n

Vocaloid (vocaloid):

not quite, remember that we leave n alone since its a variable one more time n-2+2=?

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

so basicly we just substract 2 from to then add it back correct

Vocaloid (vocaloid):

right, so what do we get in the end?

OpenStudy (anonymous):

2-2=0+2=2 right?

Vocaloid (vocaloid):

uh, not quite... let me try to explain it differently I start with n marbles. I pick up 2 marbles, then I drop 2 marbles. How many marbles do I have?

OpenStudy (anonymous):

0

Vocaloid (vocaloid):

hint: I started with n marbles

OpenStudy (anonymous):

n

Vocaloid (vocaloid):

right! so what does n-2+2=?

OpenStudy (anonymous):

n

Vocaloid (vocaloid):

good! so, on the left side we have 10-4n and on the right side we have n 10-4n = n now, let's leave 10 alone since it's not a variable what happens when we add 4n to each side?

OpenStudy (anonymous):

we get 8n

Vocaloid (vocaloid):

not quite... the left side is 10-4n, what do we get when we add 4n?

OpenStudy (anonymous):

10-4n=4n

Vocaloid (vocaloid):

not quite, please try again 10-4n+4n=?

OpenStudy (anonymous):

10-4n=6n+4n=10n?

Vocaloid (vocaloid):

not quite, remember that we leave the 10 alone since it's not a variable (doesn't have n) 10-4n+4n = ?

OpenStudy (anonymous):

oh so 10-4n+4n=10-8n

Vocaloid (vocaloid):

not quite... think of it this way: I start with 10, take away 4n, then add 4n. what's left?

OpenStudy (anonymous):

10-4n

Vocaloid (vocaloid):

not quite, let's try one more time I start with 10 marbles, then I drop 4n marbles, then I pick up 4n marbles, how many marbles do I have?

OpenStudy (anonymous):

14n? I am not sure

Vocaloid (vocaloid):

let's try something different 10-4+4=?

OpenStudy (anonymous):

14

Vocaloid (vocaloid):

hint: A-B+B = A

Vocaloid (vocaloid):

so, if A-B+B = A, then what is 10-4n+4n?

OpenStudy (anonymous):

14n

Vocaloid (vocaloid):



OpenStudy (anonymous):

10

Vocaloid (vocaloid):

great! so, the left side is 10 now let's move to the right side the right side is n, what do you get when you add 4n?

OpenStudy (anonymous):

4n

Vocaloid (vocaloid):

not quite, remember that we started with n n+4n=?

OpenStudy (anonymous):

oh so we just add 4 to the N

Vocaloid (vocaloid):

not quite... hint: n+4n = (1+4)n

OpenStudy (anonymous):

oh so the thing is 5n

Vocaloid (vocaloid):

great! left side is 10 right side is 5n 10 = 5n can you try solving for n now?

OpenStudy (anonymous):

I believe so

OpenStudy (anonymous):

I think it would equal 0,2

Vocaloid (vocaloid):

10 = 5n n = ?

OpenStudy (anonymous):

2

Vocaloid (vocaloid):

great! n = 2 going back to the original problem, m = n − 2 so what is m?

OpenStudy (anonymous):

0

Vocaloid (vocaloid):

great! m=0 n=2 so what is (m,n)?

OpenStudy (anonymous):

0,2

Vocaloid (vocaloid):

great, we're done!

OpenStudy (anonymous):

yes